CS 420 Observation Analysis Report for Spring 2011 Semester

Matthew Williamson LDCSEE, West Virginia University, Morgantown, WV {mwilli65@mix.wvu.edu}

Abstract

In this paper, we examine the issues uncovered in the CS 420 class for the Spring 2011 semester. These observations come from the author of this paper and the author alone. While we do raise some concerns about the class, we hope that the feedback provided will help improve the quality of this class and other classes within the department.

1 Introduction

In this paper, we report several observations from the Spring 2011 section of CS 420: Advanced Algorithms. The purpose of this report is to identify issues and provide suggestions on how to prevent these issues from recurring. Specifically, we will examine the following issues:

(a) Prerequisites and background information,

(b) Class participation, and

(c) Class attendance.

The remarks from this paper are based on the opinion of the author and the author alone.

2 Prerequisites and Background Information

As stated in the course catalog, the prerequisite of CS 420 is CS 221 (Analysis of Algorithms). While most of the students have taken CS 221, there were two students who were taking the prerequisite simultaneously with this course. As a result, these two students did not posses the background information in order to cover as much material as desired.

This struggle was reflected early in the course, as there were some issues understanding recurrence relations, which was a fundamental concept in CS 221. While the students claimed to have understood how to do recurrence relations, when ask to complete a few exercises, they had trouble answering the exercises.

The problems related to background information went beyond CS 221. As we delve into probabilistic analysis, permutations, and combinations, most of the students initially struggled with the concepts. This was rather surprising, considering that this material was discussed in CS 220, which was a prerequisite for CS 221.

Before we were able to go into the content intended for the course, the instructor had to go over material covered in CS 221 to make sure everyone had adequate background knowledge. However, because we had to allocate more time reviewing material that should have already been covered in previous classes, less time was spent going over supplemental material that would be both beneficial and interesting to the students. We are not implying that the material actually covered in class was not beneficial or interesting, but going over new concepts would have been nice.

The question that remains is how to prevent this issue from happening again. If the students have taken the prerequisite courses, then the only things that can be addresses are:

(a) Make sure the necessary material is covered in the prerequisite courses, and

(b) Make sure the students retain the knowledge acquired from the prerequisite courses.

However, considering that two of the students were concurrently taking CS 221 along with CS 240, there must be a problem with the course registration system. A student should not be able to register for a course without completing the prerequisite courses, unless permission is given to the student by the instructor. It is our understanding that the instructor did not give permission to any student regarding taking CS 420 without completing the prerequisite courses. Therefore, it appears there exists a flaw in the online course registration system regarding prerequisite courses.

3 Participation

4 Attendance

5 Conclusion